the width of a triangle is 15 cm less than the width. the perimeter is 98 cm. find the rectangle's dimensions.
Answer by rfer(16322)

2(x-15)+2x=98
2x-30+2x=98
4x=98+30
x=128/4
x=32
x-15=17